OpenStudy (omarbirjas):

Will medal Let f(x)=((x^2)+3x-4) and g(x)=(x+4) Find f • g and state the domain. Find f/g and state the domain. Please show work.

OpenStudy (omarbirjas):

\[f(x)=(x ^{2}+3x-4)\]

OpenStudy (omarbirjas):

^^^^ Easier way to read it.

OpenStudy (anonymous):

ey there, welcome to OpenStudy. If you have any questions about the site feel free to shoot me a message. And when asking a question, you can edit the original blue area post. Now on to your question. f(x)=(x2+3x−4) g(x)=(x+4) Part (a): f(x)×g(x) You have any thoughts on how to set about this part?

OpenStudy (omarbirjas):

@freegirl112 Not at all, all i know that f • g is just like saying f(g(x))...i think. Or something along those lines.

OpenStudy (anonymous):

That is what I just eplained

OpenStudy (omarbirjas):

Yea, that's all I know. About this....i knew that before you told me btw, just saying.

OpenStudy (anonymous):

I know that

OpenStudy (omarbirjas):

Ok now what am I supposed to do?